P5-3A Presented here are selected transactions for Norlan Inc. during September of the current year. Norlan uses a perpetual inventory system.

Sept.   2

Purchased equipment on account for $65,000, terms n/30, FOB destination.

3

Freight charges of $950 were paid by the appropriate party on the September 2 purchase of equipment.

4

Purchased supplies for $4,000 cash.

6

Purchased inventory on account from Hillary Corp. at a cost of $65,000, terms 1/15, n/30, FOB shipping point.

7

Freight charges of $1,600 were paid by the appropriate party on the September 6 inventory purchase.

8

Returned damaged goods costing $5,000 that were originally purchased from Hillary on September 6. Received a credit on account.

9

Sold goods costing $15,000 to Fischer Limited for $20,000 on account, terms 2/10, n/30, FOB destination.

10

Freight charges of $375 were paid by the appropriate party on the September 9 sale of inventory.

17

Received the balance due from Fischer.

20

Paid Hillary the balance due.

21

Purchased inventory for $6,000 cash.

22

Sold inventory costing $20,000 to Kun-Tai Inc. for $27,000 on account, terms n/30, FOB shipping point.

23

Freight charges of $500 were paid by the appropriate party on the September 22 sale of inventory.

28

Kun-Tai returned goods sold for $10,000 that cost $7,500. The merchandise was restored to inventory.

P5-12A Data for Norlan Inc. are presented in P5–3A.

Instructions

(a) Record the September transactions on Norlan’s books, assuming it uses a periodic inventory system instead of a perpetual inventory system.

(b) Assume that Norlan did not take advantage of the 1% purchase discount offered by Hillary Corp. and paid Hillary on October 3 instead of September 20. Record the entry that Norlan would make on October 3 and determine the cost of missing this purchase discount to Norlan.
Answer #1

P5-3A

Date Account Titles Debit Credit
Sep-02 Equipment $         65,000
      Accounts Payable $       65,000
(Equipment purchased on account)
Sep-03 Equipment $               950
      Cash $             950
(Payment of freight on equipment)
Sep-04 Supplies $           4,000
      Cash $         4,000
(Supplies purchased for cash)
Sep-06 Inventory $         65,000
     Accounts Payable $       65,000
(Payment of inventory on account)
Sep-07 Inventory $           1,600
     Cash $         1,600
(Paid freight in cash)
Sep-08 Accounts Payable $           5,000
      Inventory $         5,000
(Returnd damaged goods)
Sep-09 Accounts Receivable $         20,000
      Sales Revenue $       20,000
(Sales made on account)
Cost of Goods Sold $         15,000
      Inventory $       15,000
(Cost of goods sold recorded)
Sep-10 Delivery Expense $               375
      Cash $             375
(Paid freight on FOB Destination sale)
Sep-17 Cash $         19,600 =20000*98%
Sales Discount $               400 =20000*2%
       Accounts receivable $       20,000
(Cash collected net of 2% discount)
Sep-20 Accounts Payable $         60,000 =65000-5000
      Cash $       59,400 =60000*99%
      Inventory $             600 =60000*1%
(Paid net of 1% discount)
Sep-21 Inventory $           6,000
      Cash $         6,000
(Purchased inventory for cash)
Sep-22 Accounts Receivable $         27,000
      Sales Revenue $       27,000
(Sales made on account)
Cost of Goods Sold $         20,000
      Inventory $       20,000
(Cost of goods sold recorded)
Sep-23 No entry
Sep-28 Sales Returns and allowances $         10,000
      Accounts receivable $       10,000
(Sales returns)
Inventory $           7,500
      Cost of Goods Sold $         7,500
(Inventory added back)

P5-12A
(a)

Date Account Titles Debit Credit
Sep-02 Equipment $         65,000
      Accounts Payable $       65,000
(Equipment purchased on account)
Sep-03 Equipment $               950
      Cash $             950
(Payment of freight on equipment)
Sep-04 Supplies $           4,000
      Cash $         4,000
(Supplies purchased for cash)
Sep-06 Purchases $         65,000
     Accounts Payable $       65,000
(Payment of inventory on account)
Sep-07 Freight In $           1,600
     Cash $         1,600
(Paid freight in cash)
Sep-08 Accounts Payable $           5,000
      Purchase Returns and allowances $         5,000
(Returned damaged goods)
Sep-09 Accounts Receivable $         20,000
      Sales Revenue $       20,000
(Sales made on account)
Sep-10 Delivery Expense $               375
      Cash $             375
(Paid freight on FOB Destination sale)
Sep-17 Cash $         19,600 =20000*98%
Sales Discount $               400 =20000*2%
       Accounts receivable $       20,000
(Cash collected net of 2% discount)
Sep-20 Accounts Payable $         60,000 =65000-5000
      Cash $       59,400 =60000*99%
      Purchase Discount $             600 =60000*1%
(Paid net of 1% discount)
Sep-21 Purchases $           6,000
      Cash $         6,000
(Purchased inventory for cash)
Sep-22 Accounts Receivable $         27,000
      Sales Revenue $       27,000
(Sales made on account)
Sep-23 No entry
Sep-28 Sales Returns and allowances $         10,000
      Accounts receivable $       10,000
(Sales returns)

(b)

Date Account Titles Debit Credit
Oct-03 Accounts Payable $         60,000
      Cash $       60,000


Cost of missing purchase discount = $60000 x 1% = $600